OpenStudy (fairygirl0404):

I have five table to do but need help getting started please. x 1 2 3 y 2 4 8 Which table will have the greatest value of y when x = 7?

Nnesha (nnesha):

`which` I can see the only table ? where are the others?

OpenStudy (fairygirl0404):

I wasn't wanting the answer I just didn't know how to get started if x=7...what do I do? Do I multiple 2 * 7 for x? the other tables are x 1 2 3 y 14 26 38 x 1 2 3 y 3 12 27 x 1 2 3 y 12 22 32

OpenStudy (venomblast):

for the first table, for every x, y common difference of 12 the second table : for every x y is increase an ADDITIONAL of 6 with initial 9. so 3+9=12 (when x =2) 12+15=27 (when x=3) 27+21=48 (when x=4) and so on. The 4th table has a common difference of 10. Clearly the 2nd table will be much higher when x =7

OpenStudy (phi):

one thing to always try is see if the difference between y values is the same in your first table x 1 2 3 y 2 4 8 4-2 is 2 and 8-4 is 4 so it's not "linear" It is going up faster than linear however, your second table shows x 1 2 3 y 14 26 38 26-14 is 12 and 38-26 is also 12 that means it's a line with slope 12 (change in y is 12 divided by change in x is 1) y-14 = 12(x-1) in point slope form, or y= 12x +2 in slope-intercept form

OpenStudy (phi):

Your fourth table x 1 2 3 y 12 22 32 has 22-12= 10 and 32-22 = 10 so another line, with slope 10 y-12= 10(x-1) or y= 10x +2

OpenStudy (phi):

The first table look for a constant multiple. For example x 1 2 3 y 2 4 8 4 is 2*2 and 8 is 2*4 that means it's geometric, y= 2^x when x is 7 you get y = 2^7 = 128

OpenStudy (phi):

the 3rd table x 1 2 3 y 3 12 27 is a bit tricker. notice we can "factor out a 3" for the y values: 3*1 3*4 3*9 and ignoring the 3's we have 1 4 9 no constant difference and no constant multiple. so we look for squares or cubes (if we are lucky). in this case those are squares: 1^2 2^2 3^2 so the table is x 1 2 3 y 3*1^2 3*2^2 3*3^2 or in general y= 3 x^2 when x =7, you get y = 3*49= 147

Nnesha (nnesha):

I was not about to give you direct answer. just wanted to see what type of help needed here.